Thursday, 20 August

Thursday, 20 August

Invited Lectures

\underline{14:00 – 14:45}

04 Mark L. Green, UCLA, Los Angeles, USA Higher Abel-Jacobi Maps (I: p.62) Chair: Frank Schreyer

MA 005

07 Alexander Eskin, University of Chicago, Chicago, USA Counting problems and semisimple groups (I: p.91) Chair: Peter Sarnak

MA 004

08 Sergey Ivanovich Pinchuk, Indiana University, Bloomington, USA and Chelyabinsk State Univ. of Technology, Chelyabinsk, Russia Reflection Principle in Higher Dimensions (I: p.104) Chair: Jürgen Leiterer

MA 001

10 Matei Machedon, University of Maryland, College Park, USA Fourier Analysis of Null Forms and Non-linear Wave Equations (I: p.124) Chair: Michael Struwe

H 1012

12 Ruth J. Williams, University of California at San Diego, La Jolla, USA Reflecting Diffusions and Queueing Networks (I: p.152) Chair: Ludwig Arnold

H 1028

17 Alexander Schrijver, CWI and University of Amsterdam, Amsterdam, Netherlands Routing and timetabling by topological search (I: p.194) Chair: Günter M. Ziegler

H 2032

\underline{15:00 – 15:45}

04 Maurizio Cornalba, Università di Pavia, Pavia, Italy Cohomology of moduli spaces of stable curves (I: p.61) Chair: Alberto Conte

MA 005

07 Joseph Bernstein, Tel Aviv University, Tel Aviv, Israel Analytic structures on representation spaces of reductive groups (I: p.88) Chair: Wolfgang Soergel

MA 004

08 Nigel David Higson, Pennsylvania State University, University Park, USA The Baum-Connes Conjecture (I: p.98) Chair: Joachim Cuntz

MA 001

10 Frank Eric Merle, Université de Cergy-Pontoise, Cergy-Pontoise, France Blow-up Phenomenon For Critical Nonlinear Schrödinger and Zakharov Equations

(I: p.124)

Chair: Elmar Schrohe

H 1012

12 Maury D. Bramson, University of Minnesota, Minneapolis, USA State Space Collapse for Queueing Networks (I: p.141) Chair: Jennifer T. Chayes

H 1028

17 Michel X. Goemans, CORE, Louvain-La-Neuve, Belgium and MIT, Cambridge, USA Semidefinite programming and combinatorial optimization (I: p.193) Chair: András Frank

H 2032

\underline{16:15 – 17:00}

!!!

Lecture by a Fields Medalist

– to be announced – Chair: N.N.

MA 005 05 Joachim Lohkamp, Universität Augsburg, Augsburg, Germany Curvature Contents of Geometric Spaces (I: p.75) Chair: Jean-Pierre Bourguignon

MA 001

06 Tomotada Ohtsuki, Tokyo Institute of Technology, Tokyo, Japan A filtration of the set of integral homology 3-spheres (I: p.82) Chair: W.B. Raymond Lickorish

MA 004

09 Krystyna M. Kuperberg, Auburn University, Auburn, USA Counterexamples to the Seifert Conjecture (I: p.117) Chair: Urs Kirchgraber

H 1012

11 Jennifer T. Chayes, Microsoft Research, Redmond, USA Birth of the Infinite Cluster: Finite-Size Scaling in Percolation (I: p.130) Chair: Jean-Dominique Deuschel

H 1028

14 Madhu Sudan, MIT, Cambridge, USA Probabilistic checking of proofs (I: p.173) Chair: Hans J. Prömel

H 2032

\underline{17:15 – 18:00}

05 Yakov Eliashberg, Stanford University, Stanford, USA Invariants in Contact Topology (I: p.70) Chair: Jean-Pierre Bourguignon

MA 001

06 Clifford Henry Taubes, Harvard University, Cambridge, USA The Geometry of the Seiberg-Witten Invariants (I: p.86) Chair: Dieter Kotschick

MA 004

09 Håkan Eliasson, Royal Institute of Technology, Stockholm, Sweden Reducibility and point spectrum for linear quasi-periodic skew-products (I: p.112) Chair: Reiner Lauterbach

H 1012

11 Detlev Buchholz, Universität Göttingen, Göttingen, Germany Scaling algebras in local relativistic quantum physics (I: p.129) Chair: Gian Michele Graf

H 1028

14 Johan Håstad, Royal Institute of Technology, Stockholm, Sweden On approximating NP-hard optimization problems (I: p.169) Chair: Hans J. Prömel

H 2032

16 Graeme Walter Milton, The University of Utah, Salt Lake City, USA Exact relations for composites: towards a complete solution (I: p.191) Chair: Krzysztof Wilmanski MA 005